# InternVLA-A1: Unifying Understanding, Generation and Action for Robotic Manipulation

### v1

@@ Abstract
Prevalent Vision-Language-Action (VLA) models are typically built upon Multimodal Large Language Models (MLLMs) and demonstrate exceptional proficiency in semantic understanding, but they inherently lack the capability to deduce physical world dynamics. Consequently, recent approaches have shifted toward World Models, typically formulated via video prediction; however, these methods often suffer from a lack of semantic grounding and exhibit brittleness when handling prediction errors. To synergize semantic understanding with dynamic predictive capabilities, we present InternVLA-A1. This model employs a unified Mixture-of-Transformers architecture, coordinating three experts for scene understanding, visual foresight generation, and action execution. These components interact seamlessly through a unified masked self-attention mechanism. Building upon InternVL3 and Qwen3-VL, we instantiate InternVLA-A1 at 2B and 3B parameter scales. We pre-train these models on hybrid synthetic-real datasets spanning InternData-A1 and Agibot-World, covering over 533M frames. This hybrid training strategy effectively harnesses the diversity of synthetic simulation data while minimizing the sim-to-real gap. We evaluated InternVLA-A1 across 12 real-world robotic tasks and simulation benchmark. It significantly outperforms leading models like and GR00T N1.5, achieving a 14.5% improvement in daily tasks and a 40%73.3% boost in dynamic settings, such as conveyor belt sorting.
Data:InternData-A1 | Code:InternVLA-A1 | Model:InternVLA-A1 | Homepage
@@ 1 Introduction
The pursuit of intelligent generalist robots remains a cornerstone of robotics research (bu2024robodual; clark2025rad; huang2023voxposer; cui2025openhelix; fang2023anygrasp; qu2025eo1; huang2025otter; li2025hamster). To realize generalist policies, the research community favors the end-to-end learning paradigm and the Vision-Language-Action (VLA) architecture (chen2025internvlam1; black2024pi_0; pi_05; bjorck2025gr00t_n1; grootn1.5; li2025cronusvla; yang2025instructvla; zhao2025cotvla; yang2025gripper; cen2025worldvla; cheang2025gr3; lin2025onetwovla; zheng2025xvla; zhai2025walloss). Built upon Multimodal Large Language Models (MLLMs) and trained on massive real-world robot demonstration datasets, VLA models exhibit remarkable performance in daily tasks such as clothes folding and table bussing. Nevertheless, the generalizability of these models still falls short of practical application requirements. Their adaptability to scene variations remains inadequate, such as dynamic settings involving industrial conveyor belts.
@@
The generalization bottlenecks in existing policies stem from two primary issues: deficient physical world cognition and a lack of adaptive manipulation capabilities. Addressing the first challenge requires integrating foundation models, such as Multimodal Large Language Models (MLLMs) (alayrac2022flamingo; beyer2024paligemma; chen2025eagle25; touvron2023llama2) or world models (zheng2024opensora; blattmann2023svd; hacohen2024ltx; assran2025vjepa2), to enhance cognitive capacity, while addressing the latter necessitates large-scale, diverse robot datasets for skill learning. Consequently, achieving a generalist policy requires a synergistic strategy that advances both model architecture and training data.
@@
Regarding model architecture, prevalent generalist policies, such as (black2024pi_0) and GR00T N1 (bjorck2025gr00t_n1), are built upon MLLMs that map visual data into a text-based feature space. Although this grants them strong semantic understanding, text tokens are ill-suited for modeling physical laws, resulting in a deficiency in physical dynamics reasoning. Consequently, these policies are optimized for reactive perception-to-action mapping, rather than reasoning about how states will evolve under motion and contact. This limitation becomes particularly evident in dynamic environments, such as industrial conveyor settings, where understanding momentum, inertia, and contact dynamics is critical. Recent efforts attempt to incorporate foresight via World Models, notably through video prediction paradigms like VPP (hu2024vpp) and Genie Envisioner (liao2025genie_envisioner). These methods generate anticipated observations to guide decision-making, yet they often suffer from weak semantic grounding and are sensitive to prediction errors. Consequently, developing a unified architecture that tightly couples semantic understanding with robust predictive dynamics remains crucial for scalable and reliable dynamics-aware manipulation.
@@
Regarding training data, existing VLA models have gained adaptability by scaling up large real-robot datasets, and current generalist policies rely heavily on such real-robot data (walke2023bridgedata; bu2025agibotworld; wu2024robomind). For example, pioneering works collected 130,000 episodes to train RT-1 (brohan2022rt1) and RT-2 (zitkovich2023rt2), and subsequent initiatives aggregated over one million demonstrations from 22 heterogeneous robots to build Open X-Embodiment (Neill2024oxe). However, relying purely on real-world collection remains challenging: covering long-tail scene variations at scale is costly and inefficient, and high heterogeneity across embodiments further complicates joint training. For instance, even though was trained with 10,000 hours across 68 tasks from seven robot morphologies, achieving strong dexterous manipulation, its robustness against scene variations is still limited. Ultimately, end-to-end training at this scale incurs exceedingly high costs while delivering diminishing returns in long-tail scene generalization, making further scaling prohibitively expensive. In contrast, simulation presents a promising complementary approach. Its extensive libraries of scenes and objects enrich sample diversity, and domain randomization simulates scene variations to improve the policy’s robustness in changing environments. Additionally, simulation ensures more controllable trajectories by eliminating noisy data. Our prior work, InternData-A1, validated that large-scale and high-fidelity simulation data can effectively support the pre-training of VLA models. Nevertheless, simulation data suffers from the inevitable sim-to-real gap, especially in contact-rich dynamics. Therefore, synergizing the diversity of simulation with the physical fidelity of real-world data presents a promising avenue to overcome these respective limitations.
@@
To address the above generalization challenges, particularly robustness against dynamic scene variations, we propose InternVLA-A1. As depicted in Figure˜1, our model features a novel architecture integrating understanding, imagination, and action. InternVLA-A1 combines the semantic reasoning of MLLMs with the prediction capability of a World Model-style imagination module. This design effectively bridges the gap between semantics and physical dynamics, facilitating foresight-aware action generation. Furthermore, to provide the diversity needed for robustness under scene variations, we build a Data Pyramid that composes complementary data sources: large-scale open-source robot data for broad scene understanding and physical priors, simulated synthetic data for scalable and controllable coverage of diverse scenes, and a small amount of specialized real-robot data for calibrating action execution in the real world. By incorporating simulation into training as a scalable source of variation, we can expand coverage over long-tail scene configurations at low cost. At the same time, we avoid the pitfalls of simulation-only learning by anchoring the policy with real-robot data, which helps correct transfer discrepancies arising from real-world sensing and physical interactions. We validate InternVLA-A1 through extensive experiments on real-world tasks and simulation benchmarks, and observe consistent improvements over strong VLA baselines. Specifically, InternVLA-A1 outperforms baselines like and GR00T N1.5 by a significant margin, showing a 14.5% increase in success rates for daily tasks and dominant performance in dynamic tasks, with improvements of 40% for Express Sorting and 73.3% for In-motion Ingredient Picking.
2 Related work
@@
In Vision-Language-Action models, the field has shifted towards integrating the multimodal capabilities of foundation models with robotic control. RT-2 (zitkovich2023rt2) pioneered the co-finetuning of MLLMs on both robotic trajectories and Internet-scale data, enabling emergent reasoning capabilities. Diverging from RT-2’s discrete token-based approach, (black2024pi_0) combines a pre-trained MLLM with flow matching to leverage Internet-scale knowledge for controlling diverse robot embodiments. Building on this, (pi_05) integrates pre-training objectives with action prediction, thereby enhancing semantic inference for long-horizon tasks. To better synergize MLLM capabilities with real-time control, GR00T N1 (bjorck2025gr00t_n1) proposes a dual-system architecture that decouples reasoning from diffusion-based action generation. Its successor, GR00T N1.5, further improves instruction following and cross-embodiment compatibility. Distinct from these approaches, InternVLA-A1 unifies the semantic understanding of MLLMs with the dynamics‑prediction capabilities, effectively bridging the semantic-dynamic gap prevalent in existing VLA architectures.
@@
In Video prediction and World models, extensive research has focused on simulating real-world dynamics conditioned on actions to facilitate robotic control (zhang2025dreamvla; hung2025nora15; zhu2025wmpo). UniPi (Unipi) trains a text-conditioned video generator paired with an inverse dynamics model to derive actions, whereas UniSim (UniSim) employs generative modeling to construct a universal simulator for training both high-level and low-level policies. A growing body of video prediction policies—including CLOVER (bu2024clover), VPP (hu2024vpp), GR-1 (GR1), GR-2 (cheang2024gr2) and (lv2025f1)—integrates future state or video generation with action prediction, aiming to leverage the rich representations learned from large-scale video pre-training. To bridge the gap between action and visual spaces, Seer (tian2025seer) introduces an end-to-end inverse dynamics model that generates actions based on forecasted visual states. Meanwhile, the Genie Envisioner (liao2025genie_envisioner) enhances spatial understanding by utilizing a multi-view video diffusion model. Despite these advancements, these policies are often sensitive to video generation quality and lack the semantic reasoning capabilities inherent in MLLMs. In contrast, InternVLA-A1 integrates the MLLM with future state prediction, thereby internalizing the causal relationship between actions and world dynamics.
In Robotic manipulation data, datasets fall into real‑world and simulated categories. Real-world datasets (walke2023bridgedata; droid; bu2025agibotworld; wu2024robomind) capture the intrinsic physical complexity of actual environments, serving as a cornerstone for current advancements by allowing models to learn behaviors end-to-end. For instance, the RT-1 and RT-2 training pipelines were built on approximately 130,000 teleoperated episodes (brohan2022rt1; zitkovich2023rt2). The Open X-Embodiment dataset (Neill2024oxe) aggregates one million episodes from multiple sources. Nevertheless, the significant heterogeneity inherent in this data poses new challenges for training generalizable VLA models. Simulated data (gu2023maniskill2; gao2025genmanip; robotwin2.0; huang2025roboground; james2020rlbench; mees2022calvin; hua2024gensim2; im; mandlekar2023mimicgen) leverages photorealistic rendering, procedural scene generation, and domain randomization to scale up data collection efficiently. For example, Robocasa (robocasa2024) produces demonstrations by combining human teleoperation with trajectory augmentation, yet the reliance on teleoperation still demands substantial manual effort. Alternatively, GraspVLA (GraspVLA) generates action data through an automated pipeline, though its scope is currently constrained to grasping tasks. The recently proposed InternData-A1 (tian2025interndata_a1) introduces a large-scale simulated demonstration dataset containing over 630k trajectories and 7,433 hours of data across diverse embodiments, skills, tasks, and scenes. Notably, models pre-trained on InternData-A1 have demonstrated performance comparable to . In this work, we formulate a data recipe mixing InternData-A1 with real-world datasets, demonstrating that this combination enhances data diversity and effectively closes the sim-to-real gap.
@@ 3.1 Architecture Overview
The Mixture-of-Transformers (MoT) architecture, recently widely adopted in unified multimodal large language models (deng2025bagle), demonstrates strong performance across both understanding and generation tasks. Drawing inspiration from these unified paradigms,InternVLA-A1 adopts the MoT architecture to seamlessly integrate scene understanding, visual foresight, and action execution within a single framework.
As illustrated in Figure˜2, InternVLA-A1 employs three experts in a unified pipeline. The understanding expert first processes multimodal inputs to capture the environmental context. These representations then inform the generation expert, which simulates the task’s evolution by predicting future states. Finally, the action expert combines these predictive dynamics with the semantic context, utilizing Flow Matching to produce precise robot control commands.
3.2 Unified Understanding-Generation-Action Transformer
@@
Understanding Expert. The understanding expert directly adopts the architecture of existing MLLMs. In this implementation, we employ InternVL3 or Qwen3-VL, distinguished by their native multimodal capabilities and strong alignment between language and vision. We adhere to the processing pipeline of the base MLLMs: the multi-view observation at time , denoted , is encoded into visual tokens via the integrated vision encoder, while language instructions are converted into text tokens using the text tokenizer. These visual and text tokens are subsequently concatenated to form prefix tokens that condition downstream experts.
Generation Expert. Inspired by the unified multimodal model Janus Pro (chen2025januspro), we adopt a decoupled visual encoding strategy to address the divergent requirements of understanding and generation. While understanding tasks demand high-level semantic abstraction typically captured by ViT-based encoders, generation tasks require preserving fine-grained spatial structure and pixel-level fidelity. To bridge this gap, our generation expert employs a VAE-based tokenizer, widely employed in image and video generation for its ability to compress visual data into a latent space optimized for high-quality reconstruction. Specifically, we utilize the Cosmos CI continuous VAE tokenizer (agarwal2025cosmos) to encode input images into continuous latent features, providing a dense and structurally rich representation for the subsequent generative processes. The tokenizer comprises an encoder and a decoder; we denote its encoder as and use the decoder to reconstruct images from latents. Let denote the Cosmos latent at timestamp ; conditioned on and , the generation expert predicts .
@@
While generative models have advanced significantly, applying them to the VLA domain presents a significant challenge: the need for high-frequency real-time inference. Traditional image and video generation models—whether based on diffusion (rombach2022high; blattmann2023stable) or next-token prediction (sun2024Llamagen)—are typically too slow for this purpose. Even optimized solutions like SANA-Sprint (chen2025sana) require 0.16 seconds per generation, restricting control to less than 6Hz. This limitation hinders fluid robotic movement, highlighting the urgent need for faster visual generation mechanisms in embodied AI. In our preceding work (lv2025f1),we adopted a next-resolution paradigm for generating visual foresight. Although effective, this approach required iterative forward passes, which compromised the real-time inference capabilities essential for VLA tasks. Consequently, we have pursued a more aggressive design strategy to optimize performance.
We process images from three views across two timestamps (the current frame and one historical frame) using the Cosmos CI continuous VAE image tokenizer. With input images resized to , each is initially encoded into a latent feature map. Directly feeding these latent tokens into the generation expert would result in an excessive sequence length, hindering both inference efficiency and training convergence. To mitigate this, we apply a convolutional layer to compress the spatial dimensions of each latent feature to (representing each image with just 16 tokens) and use a projector to align them with the transformer’s hidden dimension. These compressed tokens are then input into the generation expert, where they attend to the prefix tokens (cached as K/V at inference) via multiple layers of masked self-attention. Their hidden states are pooled over time to form the predicted future latent , supervised against the Cosmos encoding of frame . Finally, the predicted future latent are upsampled back to the original VAE latent dimensions using a deconvolutional layer and a projector. These restored features are then fed into the Cosmos decoder to reconstruct the predicted images.
@@
Action Expert. Conditioned on the language goal , current observation (via prefix tokens ), proprioception , and the generation expert’s predicted latent tokens , the action expert predicts a target action chunk . We adopt a flow matching objective to train the VLA model.
@@
Attention Mechanism. We implement a blockwise attention mask over the concatenated token streams of the understanding, generation, and action experts. A cumulative segment mask enforces a strict information flow understanding generation action: tokens in a later block can attend to all earlier blocks, while earlier blocks cannot attend forward. The prefix block (vision + language) is fully bidirectional. The generation block is also fully bidirectional and only receives Cosmos latent tokens from frames and . The action block is split into a state token and action tokens: the state token attends only to itself and earlier blocks, while action tokens attend to the state and to each other.

### v2

@@ Abstract
Prevalent Vision-Language-Action (VLA) models are typically built upon Multimodal Large Language Models (MLLMs) and demonstrate exceptional proficiency in semantic understanding, but they inherently lack the capability to deduce physical world dynamics. Consequently, recent approaches have shifted toward World Models, typically formulated via video prediction; however, these methods often suffer from a lack of semantic grounding and exhibit brittleness in the presence of video prediction errors. To synergize semantic understanding with dynamic predictive capabilities, we present InternVLA-A1. This model employs a unified Mixture-of-Transformers architecture, coordinating three experts for scene understanding, visual foresight generation, and action execution. These components interact seamlessly through a unified masked self-attention mechanism. Building upon InternVL3 and Qwen3-VL, we instantiate InternVLA-A1 at 2B and 3B parameter scales. We pre-train these models on heterogeneous data sources over real-world robot data, synthetic simulation data, and human videos, covering over 692M frames. This hybrid training strategy effectively harnesses the diversity of synthetic simulation data while minimizing the sim-to-real gap. We evaluated InternVLA-A1 on 12 real-world robotic tasks and a simulation benchmark. The results show that InternVLA-A1 consistently outperforms prior leading models: compared with , it achieves +4.4% on static manipulation tasks and +2.6% on the RoboTwin 2.0 simulation benchmark, and delivers a +26.7% boost on dynamic manipulation tasks.
Homepage | Code:InternVLA-A1 | Model:InternVLA-A1 | Data:InternData-A1
@@ 1 Introduction
The pursuit of intelligent generalist robots remains a cornerstone of robotics research [bu2024robodual, clark2025rad, huang2023voxposer, cui2025openhelix, fang2023anygrasp, qu2025eo1, huang2025otter, li2025hamster]. Currently, the community favors the end-to-end learning paradigm and the Vision-Language-Action (VLA) architecture [chen2025internvlam1, black2024pi_0, pi_05, bjorck2025gr00t_n1, grootn1.5, li2025cronusvla, yang2025instructvla, zhao2025cotvla, yang2025gripper, cen2025worldvla, cheang2025gr3, lin2025onetwovla, zheng2025xvla, zhai2025walloss] to realize such generalist policies. Built upon Multimodal Large Language Models (MLLMs) and trained on massive real-world robot demonstrations, VLA models exhibit remarkable performance in daily tasks such as clothes folding and table bussing. Nevertheless, the generalizability of these models still falls short of practical application requirements. Their adaptability to scene variations remains inadequate, such as dynamic settings involving industrial conveyor belts.
@@
The generalization bottlenecks in existing policies stem from two primary issues: deficient physical world cognition and a lack of adaptive manipulation capabilities. Addressing the first challenge requires integrating foundation models, such as Multimodal Large Language Models (MLLMs) [alayrac2022flamingo, beyer2024paligemma, chen2025eagle25, touvron2023llama2] or world models [zheng2024opensora, blattmann2023svd, hacohen2024ltx, assran2025vjepa2], to enhance cognitive capacity, while addressing the latter necessitates large-scale and diverse robot action datasets for skill learning. Consequently, achieving generalist policies requires a synergistic strategy that advances both model architecture and training data.
@@
Regarding model architecture, prevalent generalist policies, such as [black2024pi_0], [pi_05], and GR00T N1/N1.5 [bjorck2025gr00t_n1], are built upon MLLMs that map visual data into a text-based feature space. Although this grants them strong semantic understanding, text tokens are ill-suited for modeling physical laws, resulting in a deficiency in physical dynamics reasoning. Consequently, these policies are optimized for reactive perception-to-action mapping, rather than reasoning about how states will evolve under motion and contact. This limitation becomes particularly evident in dynamic environments, such as industrial conveyor settings, where understanding momentum, inertia, and contact dynamics is critical. Recent efforts attempt to incorporate foresight via World Models, notably through video prediction paradigms like VPP [hu2024vpp] and Genie Envisioner [liao2025genie_envisioner]. These methods generate anticipated observations to guide decision-making, yet they often suffer from weak semantic grounding and are sensitive to video prediction errors. Consequently, developing a unified architecture that tightly couples semantic understanding with robust predictive dynamics is crucial for reliable dynamics-aware manipulation.
@@
Regarding training data, existing VLA models have gained adaptability by scaling up large real-robot datasets, and current generalist policies rely heavily on such real-robot data [walke2023bridgedata, bu2025agibotworld, wu2024robomind]. For example, pioneering works collected 130K episodes to train RT-1 [brohan2022rt1] and RT-2 [zitkovich2023rt2], and subsequent initiatives aggregated over one million demonstrations from 22 heterogeneous robots to build Open X-Embodiment [Neill2024oxe]. However, relying solely on real-world data collection remains challenging. Although was trained with 10,000 hours demonstrations, spanning 68 tasks from seven robot morphologies and achieving strong dexterous manipulation, it still struggles to adapt to scene variations, especially in highly dynamic environments. Further expanding real-robot datasets and covering long-tail scene variations at scale is costly and inefficient. In contrast, simulation presents a promising complementary approach. Its extensive libraries of scenes and objects enrich sample diversity, and domain randomization simulates scene variations to improve the policy’s robustness in changing environments. Additionally, simulation ensures more controllable trajectories by eliminating noisy data. Our prior work, InternData-A1 [tian2025interndata_a1], validated that large-scale and high-fidelity simulation data can effectively support the pre-training of VLA models. Nevertheless, simulation data suffers from the inevitable sim-to-real gap, especially in contact-rich dynamics. Therefore, synergizing the diversity of simulation with the physical fidelity of real-world data presents a promising avenue to overcome these respective limitations.
@@
To address the above generalization challenges, particularly robustness against dynamic scene variations, we propose InternVLA-A1. As depicted in Figure˜1, our model features a novel architecture integrating understanding, generation, and action. InternVLA-A1 combines the semantic reasoning of MLLMs with the prediction capability of a World Model-style imagination module, effectively bridging the gap between semantics and physical dynamics to facilitate foresight-aware action generation. Furthermore, we train InternVLA-A1 on three complementary data sources: large-scale simulated robot trajectories, real-robot demonstrations, and egocentric human videos. This joint training recipe enables scalable diversity from simulation, grounds action execution with real-robot interactions to reduce the sim-to-real gap, and enriches visual representations from human videos to better capture complex manipulations. By combining these three sources in a unified training pipeline, InternVLA-A1 benefits from scalable variation without sacrificing physical fidelity, leading to improved robustness and real-world transfer. We validate InternVLA-A1 through extensive experiments on 12 real-world tasks and a simulation benchmarks, and observe consistent improvements over strong VLA baselines. Specifically, InternVLA-A1 surpasses with a +4.4% improvement on real-world static tasks and +2.6% gain on the RoboTwin 2.0 simulation benchmark, alongside a substantial +26.7% boost on real-world dynamic tasks.
2 Related Works
@@
In Vision-Language-Action models, a common practice is to integrate the multimodal capabilities of foundation models with robotic control. RT-2 [zitkovich2023rt2] and OpenVLA [kim2024openvla] use a vocabulary-replacement technique that maps text tokens in an LLM to discrete action representations, thereby leveraging the LLM’s general-purpose capabilities to facilitate emergent reasoning. Diverging from discrete token-based approaches, [black2024pi_0] adopts a Mixture-of-Transformers architecture that combines a pre-trained MLLM with an action expert, and uses flow matching to output continuous actions, enabling more precise and dexterous control. Building on this, [pi_05] combines high-level sub-task prediction with low-level action prediction, thereby enhancing its capability for long-horizon tasks and cross-scene generalization. In addition, GR00T N1/N1.5 [bjorck2025gr00t_n1] adopts an approach that is closer to a dual-system paradigm, pairing a VLM for high-level reasoning with a DiT for action generation. Distinct from these approaches, InternVLA-A1 unifies the semantic understanding of MLLMs with the dynamics‑prediction capabilities, effectively bridging the semantics-dynamics gap prevalent in existing VLA architectures.
@@
In Video prediction and world models, extensive research leverages future video prediction to facilitate robotic control [zhang2025dreamvla, hung2025nora15, zhu2025wmpo]. UniPi [Unipi] trains a text-conditioned video generator paired with an inverse dynamics model to derive actions, whereas UniSim [UniSim] employs generative modeling to construct a universal simulator for training both high-level and low-level policies. GR-1 [GR1] and GR-2 [cheang2024gr2] use future image prediction as an auxiliary task to enhance the policy’s visual representations. CLOVER [bu2024clover], Seer [tian2025seer], and [lv2025f1] leverage future image prediction to guide action generation in an inverse-dynamics-like manner. As video generation models have advanced, it has become increasingly popular to incorporate a pretrained video generation model to guide action execution, as in VPP [hu2024vpp] and Genie Envisioner [liao2025genie_envisioner]. Despite these advancements, these policies are often sensitive to video generation quality and lack the semantic reasoning capabilities inherent in MLLMs. In contrast, InternVLA-A1 integrates the MLLM with future visual state prediction, thereby providing semantically grounded guidance and enabling stronger adaptability in the presence of video prediction errors.
In training data, real-world robot data, simulated synthetic data, and human video data are widely used. Real-world datasets [walke2023bridgedata, droid, bu2025agibotworld, wu2024robomind] capture realistic physical dynamics essential for end-to-end learning. RT-1 [brohan2022rt1] and RT-2 [zitkovich2023rt2] were trained on approximately 130k real-robot demonstrations. And the Open X-Embodiment dataset [Neill2024oxe] aggregates one million episodes from multiple sources. [black2024pi_0] further scales the amount of real-robot training data to 10,000 hours. However, further expanding real-robot datasets by another order of magnitude remains prohibitively expensive and operationally inefficient. Simulation-based synthetic data [gu2023maniskill2, gao2025genmanip, robotwin2.0, huang2025roboground, james2020rlbench, mees2022calvin, hua2024gensim2, im, mandlekar2023mimicgen] can fully exploit rich scene and object assets, and use domain randomization to enrich sample diversity and cover long-tail scene variations. And importantly, It provides a low-cost, efficient route for data collection. Robocasa [robocasa2024] collects data by teleoperating robots in simulation and then performing sample augmentation. Its initial data collection still requires manual effort. GraspVLA [GraspVLA] can automatically generate robot trajectories, but its skills are limited to grasping. More recently, InternData-A1 [tian2025interndata_a1] proposes an automated robot trajectory synthesis pipeline that covers the manipulation of rigid, articulated, deformable, and fluid objects, and constructs a large-scale simulated demonstration dataset with over 630k trajectories and 7,433 hours of data spanning diverse embodiments, skills, tasks, and scenes. There is a vast amount of video data on the internet that remains underexplored for its potential to benefit the pretraining of manipulation policies. Human videos [egodex, damen2020epic, grauman2024ego4d, goyal2017something] provide rich visual priors without costly teleoperation, proving effective for human-to-robot transfer [emergence_h2r, bi2025h_rdt]. For instance, Ego4D [grauman2024ego4d] contains 3,670 hours of daily activities. Tailored for manipulation, EgoDex [egodex] curates 829 hours of egocentric dexterous manipulation footage with paired 3D hand and finger tracking. Some studies [cheang2024gr2, motus] have incorporated human video data to enrich the diversity of pre-training datasets. In this work, we formulate a data recipe mixing real-world datasets, simulated data, and human videos, demonstrating that this combination enhances data diversity and effectively closes the sim-to-real gap.
@@ 3.1 Architecture Overview
The Mixture-of-Transformers (MoT) architecture, recently widely adopted in unified multimodal large language models [deng2025bagle], demonstrates strong performance across both understanding and generation tasks. Drawing inspiration from these unified paradigms, InternVLA-A1 adopts the MoT architecture to seamlessly integrate scene understanding, visual foresight, and action execution within a single framework.
As illustrated in Figure˜2, InternVLA-A1 employs three experts in a unified pipeline. The understanding expert first processes multimodal inputs to capture the environmental context. These representations then inform the generation expert, which simulates the task’s evolution by predicting future visual states. Finally, the action expert combines these predictive dynamics with the semantic context, utilizing Flow Matching to produce precise robot control commands.
3.2 Core Components
@@
Understanding Expert. The understanding expert directly adopts the architecture of existing MLLMs. In this implementation, we employ InternVL3 or Qwen3-VL, distinguished by their native multimodal capabilities and strong alignment between language and vision. We adhere to the processing pipeline of the base MLLMs: the multi-view observation at time , denoted , is encoded into visual tokens via the integrated vision encoder, while language instructions are converted into text tokens using the text tokenizer. These visual and text tokens are subsequently processed by the transformer blocks of understanding expert, to form contextual embeddings . These embeddings serve as a shared context memory that is made accessible to downstream experts through masked self-attention, enabling the generation and action experts to attend to semantic scene context when predicting future latents and control commands.
@@
Generation Expert. While video generation models have seen substantial progress, applying them to manipulation policies presents a significant challenge due to the requirements for high-frequency real-time inference. Mainstream image or video generation architectures, whether based on diffusion [rombach2022high, blattmann2023stable] or next-token prediction [sun2024Llamagen], are typically too computationally intensive for end-to-end control. Even optimized solutions such as SANA-Sprint [chen2025sana] require 0.16 seconds per generation on one RTX 4090 GPU, restricting control frequencies to no more than 6Hz. Recent attempts [motus, dreamzero2025] leverage large-scale pretrained video foundation models and tailor them for action control; however, these methods lack sufficient real-time performance for high-dynamic scenarios. A notable example is DreamZero [dreamzero2025]: even after a 38 speedup from extensive engineering optimization, it attains only 7Hz on a GB200 GPU, presenting a significant barrier to robotic edge deployment and real-time inference. Instead, we choose to implement a lightweight generation module. In our preceding work [lv2025f1], we adopted a next-resolution paradigm for generating visual foresight. Although effective, this approach required iterative forward passes, which compromised the real-time inference capabilities essential for VLA tasks. In this work, we have pursued a more efficient and effective design strategy.
The detailed architectural design of the Generation Expert is illustrated in Figure˜3. Specifically, we expand upon three key specifics: Input Tokenization, Token Compression, and Parallel Decoding.
Input Tokenization: the Generation Expert takes as input images from three perspectives, including a head view and two wrist views. To capture temporal dynamics, we sample frames from two timestamps: the current timestep and a historical timestep . This results in a total of 6 input images, each resized to a resolution of . Inspired by the unified multimodal model Janus Pro [chen2025januspro], we adopt a decoupled visual encoding strategy to address the divergent requirements of understanding and generation. Unlike understanding tasks demanding high-level semantic abstraction typically captured by ViT-based encoders, generation tasks require preserving fine-grained spatial structure and pixel-level fidelity. Accordingly, our generation expert employs a VAE-based tokenizer, widely employed in image and video generation for its ability to compress visual data into a latent space optimized for high-quality reconstruction. Specifically, we utilize the COSMOS CI88 continuous VAE image tokenizer [agarwal2025cosmos] to encode these input images. In its raw form, each image is encoded into a latent grid, resulting in 1,024 tokens per image. Directly feeding these latent tokens into the generation expert would result in an excessive sequence length, hindering both inference efficiency and training convergence. Therefore, token compression is required.
Token Compression: to address the sequence length challenge, we implement a token compression mechanism. We apply a convolution layer with an kernel to downsample the latent grid, reducing the representation of each image to (16 tokens). Consequently, the input sequence for the 6 images (3 views 2 timestamps) is compacted to just 96 tokens. These tokens are then passed through a projector to align with the transformer blocks’ hidden dimension. Within the transformer blocks, tokens from the Generation Expert attend to both themselves and the Understanding Expert’s tokens via a unified masked self-attention mechanism, utilizing the preceding Understanding Expert tokens as the KV cache. Subsequently, the processed tokens undergo parallel decoding to generate future images.
Parallel Decoding: following the transformer processing, the output remains a sequence of 96 tokens. We apply temporal average pooling along the time axis to aggregate information from the two timestamps, resulting in 48 tokens representing the three views (16 tokens per view). These tokens are processed by a projector and subsequently upsampled via a deconvolution layer back to a grid. Finally, the COSMOS VAE decoder reconstructs the predicted future frames (corresponding to ). Notably, our approach avoids auto-regressive next-token prediction. Instead, we employ a single-forward parallel decoding strategy, generating all tokens for the future frames simultaneously. We find that this non-autoregressive approach is not only computationally efficient but also sufficient to provide effective visual guidance for the subsequent action execution.
@@
Action Expert. Conditioned on the latent features produced by the understanding and generation experts, together with the proprioceptive state , the action expert predicts a target action chunk . We adopt a flow matching objective to train the VLA model.
@@
Attention Mechanism. We implement a blockwise attention mask over the concatenated token streams of the understanding, generation, and action experts. A cumulative segment mask enforces a strict information flow understanding generation action: tokens in a later block can attend to all earlier blocks, while earlier blocks cannot attend forward. Within the transformer blocks of both the understanding and generation experts, the tokens are fully bidirectionally attended. The tokens processed by the action expert’s transformer blocks are split into a state token and action tokens. The state token attends only to itself and the tokens from earlier blocks, while action tokens attend to the state token, previous block tokens, and each other.